Chris invests a total of $31,000 in two accounts paying 15% and 7% annual interest, respectively. How much was invested in each account if, after one year, the total interest was $3,290.00.
$ was invested at 15% and
$ was invested at 7%.

Answer by macston(5194)

.
F=amount at 15%; S=amount at 7%
F+S=$31000
S=$31000-F Use this to substitute for S below.
.
0.07S+0.15F=$3290 Substitute for S from above.
0.07($31000-F)+0.15F=$3290
$2170-0.07F+0.15F=$3290
0.08F=$1120
F=$14000 ANSWER 1: $14000 was invested at 15%.
.
S=$31000-$14000
S=$17000 ANSWER 2: $17000 was invested at 7%.
.
CHECK:
.
0.07S+0.15F=$3290
0.07($17000)+0.15($14000)=$3290
$1190+$2100=$3290
$3290=$3290